OpenStudy (anonymous):

Well, we found another one:) 4/x+2 - 1/x = 2x-1/x2 (that is squared) + 2x.

OpenStudy (anonymous):

\[{4 \over x+2} - {1 \over x} = {2x -1 \over x^2 + 2x}\] So start by putting each of the fractions over a common denominator.

OpenStudy (anonymous):

Ok, I will start doing that and get back with you:) Thank you!

OpenStudy (anonymous):

Umm, should x or x\[^{x2}\] +2x be the common denominator?

OpenStudy (anonymous):

It keeps messing my numbers up. x or x2 plus 2x

OpenStudy (anonymous):

I think \[x^2 + 2x\] will be the common denominator.

OpenStudy (anonymous):

We are trying to work it, so you do not think I left:)

OpenStudy (anonymous):

No problem =)

OpenStudy (anonymous):

Is is possible to get the next step?

OpenStudy (anonymous):

The next step would be to combine the two fractions on the left side.

OpenStudy (anonymous):

Did you get the new form yet?

OpenStudy (anonymous):

Ok, we got that part down:)

OpenStudy (anonymous):

So what do you have now?

OpenStudy (anonymous):

4x2 + 8x - x2 - 2x = 2x-1

OpenStudy (anonymous):

That doesn't look quite right.

OpenStudy (anonymous):

\[{4 \over x+2} = {4x \over x(x+2)} = {4x \over x^2+2x}\] \[{-1 \over x} = {-1(x+2) \over x(x+2)} = {-x -2 \over x^2+2x}\] \[\implies {4 \over x+2} - {1 \over x} = {4x -x -2\over x^2+2x}\]

OpenStudy (anonymous):

WOW! How did you even get that to work? I will write it down and be right back.

OpenStudy (anonymous):

You do what you do for regular fractions. Figure out what factors you need to multiply by top and bottom to get the denominator you want. Since x+2 needs a factor of x to become \(x^2 + 2x\) we multiply top and bottom of that fraction by x. Since x needs a factor of x+2 to become\(x^2 + 2x\) we multiply top and bottom of that fraction by (x+2)

OpenStudy (anonymous):

I know this sounds bad, but what is the next step?

OpenStudy (anonymous):

Well you have something that looks like this now right? \[{4x -x -2\over x^2+2x} ={2x -1 \over x^2 + 2x} \] Start by combining the sum you have on the left side there.

OpenStudy (anonymous):

Ok:)

OpenStudy (anonymous):

So what do you have?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Sounds about right.

OpenStudy (anonymous):

Wait no.

OpenStudy (anonymous):

Almost tricked me.

OpenStudy (anonymous):

Oh wait. that is right. I tricked myself!

OpenStudy (anonymous):

Just goes to show. Don't do things only in your head ;)

OpenStudy (anonymous):

yea!!! God is so good, thatnk you!! I do not know how you are able to keep up with this stuff save for divine grace. You are a math expert!

OpenStudy (anonymous):

Again, happy to help. =)

OpenStudy (anonymous):

Can we give u another one?

OpenStudy (anonymous):

x/2 + 2x-1/3 = 9

OpenStudy (anonymous):

Same type deal. need a common denominator.

OpenStudy (anonymous):

is it 6?

OpenStudy (anonymous):

Oh, and is it \[{x \over 2} + {2x-1 \over 3} = 9\]?

OpenStudy (anonymous):

Yes it is!

OpenStudy (anonymous):

Parentheses help disambiguate.

OpenStudy (anonymous):

Ok then yes, common denominator is 6.

OpenStudy (anonymous):

So what would you multiply the top and bottom of the first fraction by to get a 6 in the denominator?

OpenStudy (anonymous):

3 and 2

OpenStudy (anonymous):

Yeah, 2 for the second one. You may as well also put the 9 over a 6 denominator while you're at it. Makes things easier.

OpenStudy (anonymous):

should it be 9*6?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

So what have you got now?

OpenStudy (anonymous):

ok, so if the problem turns out to be 6x+12-6=54 what will the bottom numbers be?

OpenStudy (anonymous):

Should be 6x + 12x - 6 = 54, and on bottom you will have all 6's (by design)

OpenStudy (anonymous):

So you can actually just ignore the denominators now.

OpenStudy (anonymous):

\[{6x + 12x - 6 \over 6} = {54 \over 6} \implies 6x + 12x - 6 = 54\]

OpenStudy (anonymous):

Ok, we have that part.

OpenStudy (anonymous):

So now it's just a matter of isolating your x and dividing away the coefficient.

OpenStudy (anonymous):

We tried that, but we got 18x=60. Is that right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Wait, something is wrong.

OpenStudy (anonymous):

Would you mind providing the next step? We cannot get it right.

OpenStudy (anonymous):

Well the next step should be to divide both sides by 18, but I think we have a mistake further back.

OpenStudy (anonymous):

\[{x \over 2} + {2x-1 \over 3} = 9\]\[\implies {x \over 2}(\frac{3}{3}) + {(2x-1) \over 3}(\frac{2}{2}) = 9(\frac{6}{6})\]\[\implies {3x + 2(2x - 1) \over 6} = {54 \over 6} \]\[ \implies {3x + 4x - 2 \over 6} = {54 \over 6} \]\[\implies 3x + 4x - 2 = 54\]

OpenStudy (anonymous):

That's better. Somehow I didn't notice that you'd goofed your denominator conversion.

OpenStudy (anonymous):

7x=56? x=8?

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

You have the basic stuff down. You need to practice your conversion of fractions.

OpenStudy (anonymous):

Praise God that we got the right answer! Thank you again! I think that is all. God bless

OpenStudy (anonymous):

Knowing what you did wrong is more important than getting it right. Be sure to look over your solution and see what went wrong.

OpenStudy (anonymous):

And why it's not correct.

OpenStudy (anonymous):

We will do that. Thank you again! God bless